OpenStudy (anonymous):

I will defently give a medal, maybe fan! What is the value of w? 5(9 − 3) = w • 9 – 5 • 3 w = ______

OpenStudy (uri):

*defently* okay...

OpenStudy (anonymous):

First you distribute the 5 on the left side. Remember the order of operations? Use that on the right side.

OpenStudy (uri):

Did you try doing it? @Kyliejones Please show your work...You apply the PEMDAS rule.

OpenStudy (anonymous):

45 - 15 = 9w - 15 45 = 9w Do you know how to solve from there?

OpenStudy (compassionate):

5(9 − 3) = w • 9 – 5 • 3 First distribute what you can. You see the five can be distributed to 9 and - 3 45 - 15 = w • 9 - 5 • 3 Thereafter, you need to multiply next. That is, multiply w • and -5 • 3 After that, add like terms then solve.

OpenStudy (anonymous):

I honestly dont really understand im in the 9th grade and they gave this to everyone, But i kinda still dont get it?

OpenStudy (anonymous):

w = 5

OpenStudy (anonymous):

At first i got 5.5 but i was like no thats wrong? so i didint put the answer in

OpenStudy (anonymous):

PEMDAS Please Excuse My Dear Aunt Sally Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (anonymous):

Thank you! :)

OpenStudy (compassionate):

Do you understand it yet? Kylie? Explain it to me.

OpenStudy (anonymous):

Do you remember the properties of multiplication?

OpenStudy (anonymous):

Yes i do There are four properties involving multiplication that will help make problems easier to solve. They are the commutative, associative, multiplicative identity and distributive properties. Commutative property: When two numbers are multiplied together, the product is the same regardless of the order of the multiplicands. For example 4 * 2 = 2 * 4 Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4) Multiplicative Identity Property: The product of any number and one is that number. For example 5 * 1 = 5. Distributive property: The sum of two numbers times a third number is equal to the sum of each addend times the third number. For example 4 * (6 + 3) = 4*6 + 4*3 Am i right?

OpenStudy (compassionate):

Yes, you are correct. We can pretty much ignore the distributive property, because that's just fancy word for ,"multiplication." You see, in 5(9 − 3) = w • 9 – 5 • 3 We know that 5 has to be multiplied by 9 - 3, then w * 9 and -5 * 3. We just need to multiply first, that's all. Then add like terms, then finally divide at the end to solve.

OpenStudy (anonymous):

Thats sometimes hard because i dont understand But thank you so much for you helping me! Im actully writeing it all down :)

OpenStudy (anonymous):

*writing

OpenStudy (compassionate):

Yes! It's quite simple. If you need anymore help, tag me.

OpenStudy (anonymous):

I will! :)